# Topological restrictions on relatively Anosov representations

### v1

@@ 7. Questions
We know from the above that if $(\Gamma,\operatorname{\mathcal{P}})$ is a relatively hyperbolic group and $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\mathsf{SL}_{d}(\mathbb{R})$ is a relatively $k$-Anosov representation with $d\geq 2k$, then the dimension of $\partial(\Gamma,\operatorname{\mathcal{P}})$ is bounded above by $d-k-1$ (see also [CT20] for the non-relative case). We also know that, when $k=1$ and $\partial(\Gamma,\operatorname{\mathcal{P}})$ is connected and not a circle, the image of $\rho$ preserves a domain in $\mathbb{P}(\operatorname{\mathbb{R}}^{d})$ and hence $\partial(\Gamma,\operatorname{\mathcal{P}})$ embeds into a $(d-2)$-sphere. As noted in §5.2 above, a natural question that then arises is the following:
@@
Beyond relatively Anosov representations, a larger, more flexible class of representations generalizing geometric finiteness in the setting of higher-rank Lie groups is given by the extended geometrically finite (EGF) representations introduced by Weisman [Wei22]. Roughly speaking, these are defined in terms of extended boundary maps with good dynamical properties. In general, these maps are no longer homeomorphisms, but rather equivariant surjections from subsets of the flag space to the Bowditch boundary of the domain group. Our results here would not apply directly, but it is possible that some of the techniques can be adapted to the study of EGF representations. In particular, we can ask some of the following questions:

### v2

@@ 7. Questions
We know from the above that if $(\Gamma,\operatorname{\mathcal{P}})$ is a relatively hyperbolic group and $\rho\colon(\Gamma,\operatorname{\mathcal{P}})\to\mathsf{SL}_{d}(\mathbb{R})$ is a relatively $k$-Anosov representation with $d\geq 2k$, then the dimension of $\partial(\Gamma,\operatorname{\mathcal{P}})$ is bounded above by $d-k-1$ (see also [CanaryTsouvalas] for the non-relative case). We also know that, when $k=1$ and $\partial(\Gamma,\operatorname{\mathcal{P}})$ is connected and not a circle, the image of $\rho$ preserves a domain in $\mathbb{P}(\operatorname{\mathbb{R}}^{d})$ and hence $\partial(\Gamma,\operatorname{\mathcal{P}})$ embeds into a $(d-2)$-sphere. As noted in §5.2 above, a natural question that then arises is the following:
@@
Beyond relatively Anosov representations, a larger, more flexible class of representations generalizing geometric finiteness in the setting of higher-rank Lie groups is given by the extended geometrically finite (EGF) representations introduced by Weisman [Teddy_EGF]. Roughly speaking, these are defined in terms of extended boundary maps with good dynamical properties. In general, these maps are no longer homeomorphisms, but rather equivariant surjections from subsets of the flag space to the Bowditch boundary of the domain group. Our results here would not apply directly, but it is possible that some of the techniques can be adapted to the study of EGF representations. In particular, we can ask some of the following questions:
@@
[…]
Max Planck Institute for Mathematics in the Sciences, Inselstraße 22, 04103 Leipzig, Germany
E-mail address: konstantinos.tsouvalas@mis.mpg.de
Department of Mathematics, University of Wisconsin – Madison, WI 53706, USA
E-mail address: fzhu52@wisc.edu